# Complexity reduction of physical models: an equation-free approach by means of scaling

### v1

@@ Appendix B Theorems
We consider the vectors $\vec{l}_{1},\dots,\vec{l}_{N_{\mathsf{l}}}\in\mathbb{R}^{N_{\mathsf{p}}}$ of exponents in (B.21). As explained in the proof of Theorem B.4, we have that Theorems B.2 and B.3 combined with Properties A.4, A.5 and A.6 ensure (B.22) for each vector $\vec{l}_{1},\dots,\vec{l}_{N_{\mathsf{l}}}$. This can be understood as follows. Let us denote B as the matrix whose columns are the vectors $\vec{b}_{1},\dots,\vec{b}_{N_{\mathsf{0}}}$ (B.12), i.e.,
@@
$\textbf{B}\coloneqq\begin{pmatrix}\vec{b}_{1}&\vec{b}_{2}&\cdots&\vec{b}_{N_{\mathsf{0}}}\end{pmatrix}\in\mathbb{R}^{N_{\mathsf{p}}\times N_{\mathsf{0}}},\quad\mbox{with}\quad N_{\mathsf{0}}<N_{\mathsf{p}}.$ (B.26)
@@
Each of the vectors $\vec{l}_{1},\dots,\vec{l}_{N_{\mathsf{l}}}$ in (B.22) belongs to the column space of such a matrix B (B.26), which is equivalent to the orthogonal complement of the null space (kernel) of $\textbf{B}^{T}$, i.e.,
@@
$\vec{l}_{i}\in\mathrm{colsp}(\textbf{B})\equiv\left(\mathrm{ker}(\textbf{B}^{T})\right)^{\bot},\quad\forall i=1,\dots,N_{\mathsf{l}}.$ (B.27)
@@
One has $N_{\mathsf{0}}=N_{\mathsf{p}}-\mathrm{rank}(\textbf{M})<N_{\mathsf{p}}$ in (B.26) owing to $\mathrm{rank}(\textbf{M})\geq 1$. The latter inequality follows from Property A.6 guaranteeing that M (B.8) cannot be a matrix whose entries are all equal to zero. Then, since $\mathrm{rank}(\textbf{B}^{T})=\mathrm{rank}(\textbf{B})=N_{\mathsf{0}}$ and $\mathrm{dim}\left(\mathrm{ker}(\textbf{B}^{T})\right)=N_{\mathsf{p}}-N_{\mathsf{0}}>0$, we can consider $N_{\mathsf{p}}-N_{\mathsf{0}}$ vectors $\vec{v}_{1},\dots,\vec{v}_{N_{\mathsf{p}}-N_{\mathsf{0}}}\in\mathbb{R}^{N_{\mathsf{p}}}$ to form a basis of $\mathrm{ker}(\textbf{B}^{T})$. The assertion (B.27) is equivalently stated as each of $\vec{l}_{1},\dots,\vec{l}_{N_{\mathsf{l}}}$ being perpendicular to each of the basis’ vectors $\vec{v}_{1},\dots,\vec{v}_{N_{\mathsf{p}}-N_{\mathsf{0}}}$, i.e.,
@@
$\textbf{V}^{T}\,\vec{l}_{i}=\vec{0},\quad\textbf{V}\coloneqq\begin{pmatrix}\vec{v}_{1}&\vec{v}_{2}&\cdots&\vec{v}_{N_{\mathsf{p}}-N_{\mathsf{0}}}\end{pmatrix}\in\mathbb{R}^{N_{\mathsf{p}}\times(N_{\mathsf{p}}-N_{\mathsf{0}})},\quad\forall i=1,\dots,N_{\mathsf{l}}.$ (B.28)
@@
Property A.5 guarantees that the numerical values $\kappa_{i}$ and $\theta_{j}$ of all $\tilde{\kappa}_{i}$ in (A.2) and $\tilde{\theta}_{j}$ in (A.3) are strictly positive. Then, one can plug (A.2) and (A.3) into (A.1) to get
$\lambda_{i}=\tilde{p}_{1}^{\sum_{j=1}^{N_{\mathsf{x}}}c_{i,j}\,t_{1,j}\,+\,d_{i,1}}\,\tilde{p}_{2}^{\sum_{j=1}^{N_{\mathsf{x}}}c_{i,j}\,t_{2,j}\,+\,d_{i,2}}\,\cdots\,\tilde{p}_{N_{\mathsf{p}}}^{\sum_{j=1}^{N_{\mathsf{x}}}c_{i,j}\,t_{N_{\mathsf{p}},j}\,+\,d_{i,N_{\mathsf{p}}}},\quad\forall i=1,\dots,N_{\mathsf{l}}.$ (B.29)
@@
By using (B.29) and the independence of physical parameters $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$ (Property A.4), the vectors $\vec{l}_{1},\dots,\vec{l}_{N_{\mathsf{l}}}$ of exponents in (B.21) can be computed in terms of $\vec{t}_{1},\dots,\vec{t}_{N_{\mathsf{x}}}$ (B.24):
$\vec{l}_{i}=c_{i,1}\,\vec{t}_{1}\,+\,c_{i,2}\,\vec{t}_{2}\,+\,\cdots\,+\,c_{i,N_{\mathsf{x}}}\,\vec{t}_{N_{\mathsf{x}}}\,+\,\vec{d}_{i},\quad\vec{d}_{i}\coloneqq(d_{i,1},d_{i,2},\dots,d_{i,{N_{\mathsf{p}}}})^{T},\quad\forall i=1,\dots,N_{\mathsf{l}}.$ (B.30)
Inserting $\vec{l}_{i}$ (B.30) into (B.28) gives
$c_{i,1}\,\textbf{V}^{T}\,\vec{t}_{1}\,+\,c_{i,2}\,\textbf{V}^{T}\,\vec{t}_{2}\,+\,\cdots\,+\,c_{i,N_{\mathsf{x}}}\,\textbf{V}^{T}\,\vec{t}_{N_{\mathsf{x}}}=-\textbf{V}^{T}\,\vec{d}_{i},\quad\forall i=1,\dots,N_{\mathsf{l}},$ (B.31)
that reads in terms of the vector $\vec{t}$ (B.24) of unknowns as
$\textbf{T}\,\vec{t}=\vec{\tau},\quad\textbf{T}\in\mathbb{R}^{N_{\mathsf{t}}\times(N_{\mathsf{p}}\,N_{\mathsf{x}})},\quad\vec{\tau}\in\mathbb{R}^{N_{\mathsf{t}}},\quad N_{\mathsf{t}}\coloneqq N_{\mathsf{l}}\,(N_{\mathsf{p}}-N_{\mathsf{0}}),$ (B.32)
@@
$\textbf{T}=\begin{pmatrix}c_{1,1}\,\textbf{V}^{T}&c_{1,2}\,\textbf{V}^{T}&\cdots&c_{1,N_{\mathsf{x}}}\,\textbf{V}^{T}\\ c_{2,1}\,\textbf{V}^{T}&c_{2,2}\,\textbf{V}^{T}&\cdots&c_{2,N_{\mathsf{x}}}\,\textbf{V}^{T}\\ \vdots&\vdots&\ddots&\vdots\\ c_{N_{\mathsf{l}},1}\,\textbf{V}^{T}&c_{N_{\mathsf{l}},2}\,\textbf{V}^{T}&\cdots&c_{N_{\mathsf{l}},N_{\mathsf{x}}}\,\textbf{V}^{T}\end{pmatrix},\quad\vec{\tau}=-\begin{pmatrix}\textbf{V}^{T}\,\vec{d}_{1}\\ \textbf{V}^{T}\,\vec{d}_{2}\\ \vdots\\ \textbf{V}^{T}\,\vec{d}_{N_{\mathsf{l}}}\end{pmatrix},$ (B.33)
@@
with V and $\vec{d}_{1},\dots,\vec{d}_{N_{\mathsf{l}}}$ defined as in (B.28) and (B.30), respectively. By using any solution $\vec{t}\coloneqq(\vec{t}_{1},\vec{t}_{2},\dots,\vec{t}_{N_{\mathsf{x}}})^{T}$ to (B.32)-(B.33) in (B.30), the resulting vectors $\vec{l}_{1},\dots,\vec{l}_{N_{\mathsf{l}}}$ (B.30) fulfill (B.28) that implies (B.27). In other words, such $\vec{l}_{1},\dots,\vec{l}_{N_{\mathsf{l}}}$ belong to the column space of the matrix B (B.26) generated by the basis’ vectors $\vec{b}_{1},\dots,\vec{b}_{N_{\mathsf{0}}}$ (B.12). Thus, each of the computed $\vec{l}_{1},\dots,\vec{l}_{N_{\mathsf{l}}}$ admits the unique vector $\vec{\alpha}_{i}$ (B.22) of coordinates with respect to the basis $\vec{b}_{1},\dots,\vec{b}_{N_{\mathsf{0}}}$ (B.12). That is to say, each of the $N_{\mathsf{l}}$ linear systems written in (B.22), i.e., $\textbf{B}\,\vec{\alpha}_{i}=\vec{l}_{i}$, with $i=1,\dots,N_{\mathsf{l}}$, can be solved providing the unique solution $\vec{\alpha}_{i}\in\mathbb{R}^{N_{\mathsf{0}}}$:
@@
$\vec{\alpha}_{i}=\textbf{B}^{+}\,\vec{l}_{i},\quad\textbf{B}^{+}\coloneqq\left(\textbf{B}^{T}\,\textbf{B}\,\right)^{-1}\textbf{B}^{T}\in\mathbb{R}^{N_{\mathsf{0}}\times N_{\mathsf{p}}},\quad\forall i=1,\dots,N_{\mathsf{l}},$ (B.34)
@@
where $\textbf{B}^{+}$ is the pseudo-inverse of the rectangular matrix B (B.26). The inverse matrix $\left(\textbf{B}^{T}\,\textbf{B}\,\right)^{-1}\in\mathbb{R}^{N_{\mathsf{0}}\times N_{\mathsf{0}}}$ in (B.34) is well defined because $\mathrm{rank}(\textbf{B})=N_{\mathsf{0}}<N_{\mathsf{p}}$ implies that $\mathrm{rank}(\textbf{B}^{T}\,\textbf{B})=N_{\mathsf{0}}$. By plugging $\vec{l}_{i}$ (B.30) into (B.34), it follows
$c_{i,1}\,\textbf{B}^{+}\,\vec{t}_{1}\,+\,c_{i,2}\,\textbf{B}^{+}\,\vec{t}_{2}\,+\,\cdots\,+\,c_{i,N_{\mathsf{x}}}\,\textbf{B}^{+}\,\vec{t}_{N_{\mathsf{x}}}=\vec{\alpha}_{i}\,-\,\textbf{B}^{+}\,\vec{d}_{i},\quad\forall i=1,\dots,N_{\mathsf{l}},$ (B.35)
that reads in terms of the vector $\vec{t}$ (B.24) of unknowns as
$\textbf{A}\,\vec{t}=\vec{\alpha}\,-\,\vec{\omega},\quad\textbf{A}\in\mathbb{R}^{(N_{\mathsf{l}}\,N_{\mathsf{0}})\times(N_{\mathsf{p}}\,N_{\mathsf{x}})},\quad\vec{\alpha},\vec{\omega}\in\mathbb{R}^{N_{\mathsf{l}}\,N_{\mathsf{0}}},$ (B.36)
@@
$\textbf{A}=\begin{pmatrix}c_{1,1}\,\textbf{B}^{+}&c_{1,2}\,\textbf{B}^{+}&\cdots&c_{1,N_{\mathsf{x}}}\,\textbf{B}^{+}\\ c_{2,1}\,\textbf{B}^{+}&c_{2,2}\,\textbf{B}^{+}&\cdots&c_{2,N_{\mathsf{x}}}\,\textbf{B}^{+}\\ \vdots&\vdots&\ddots&\vdots\\ c_{N_{\mathsf{l}},1}\,\textbf{B}^{+}&c_{N_{\mathsf{l}},2}\,\textbf{B}^{+}&\cdots&c_{N_{\mathsf{l}},N_{\mathsf{x}}}\,\textbf{B}^{+}\end{pmatrix},\quad\vec{\omega}=\begin{pmatrix}\textbf{B}^{+}\,\vec{d}_{1}\\ \textbf{B}^{+}\,\vec{d}_{2}\\ \vdots\\ \textbf{B}^{+}\,\vec{d}_{N_{\mathsf{l}}}\end{pmatrix},$ (B.37)
@@
where $\vec{\alpha}$ is defined as in (B.39), being the vector $\in\mathbb{R}^{N_{\mathsf{l}}\,N_{\mathsf{0}}}$ whose entries are the exponents $\alpha_{i,k}$ in (B.20). The matrix $\textbf{B}^{+}$ and the vectors $\vec{d}_{1},\dots,\vec{d}_{N_{\mathsf{l}}}$ are defined as in (B.34) and (B.30), respectively. Both linear systems (B.32)-(B.33) and (B.36)-(B.37) must be satisfied by the vector $\vec{t}$ (B.24) of unknowns, yielding to the matrix equation $\textbf{Y}\,\vec{t}=\vec{\gamma}$ (B.25) for $\vec{t}$, with Y and $\vec{\gamma}$ given by
@@
$\textbf{Y}=\begin{pmatrix}\textbf{A}\\ \textbf{T}\end{pmatrix}\in\mathbb{R}^{(N_{\mathsf{l}}\,N_{\mathsf{p}})\times(N_{\mathsf{p}}\,N_{\mathsf{x}})},\quad\vec{\gamma}=\begin{pmatrix}\vec{\alpha}-\vec{\omega}\\ \vec{\tau}\end{pmatrix}\in\mathbb{R}^{N_{\mathsf{l}}\,N_{\mathsf{p}}}.$ (B.38)
@@
The entries of Y (B.38) are explicitly known real numbers, as given by those of A (B.37) and T (B.33). Since the components of $\vec{\omega}$ (B.37) and $\vec{\tau}$ (B.33) are explicitly known, the entries of $\vec{\gamma}$ (B.38) are explicit affine functions of the exponents $\alpha_{i,k}$ in (B.20), collected by the vector $\vec{\alpha}$ (B.39).
Theorem B.6 provides a way for finding unknown exponents $\alpha_{i,k}$ in (B.20) (Theorem B.4).
Theorem B.6.
Given Properties A.1 to A.6 and the consequent expression (B.20) of coefficients $\lambda_{1},\dots,\lambda_{N_{\mathsf{l}}}$ (13), the vector
$\vec{\alpha}\coloneqq\begin{pmatrix}\vec{\alpha}_{1}\\ \vec{\alpha}_{2}\\ \vdots\\ \vec{\alpha}_{N_{\mathsf{l}}}\end{pmatrix}\in\mathbb{R}^{N_{\mathsf{l}}\,N_{\mathsf{0}}},\quad\mbox{with}\quad\vec{\alpha}_{i}\coloneqq\begin{pmatrix}\alpha_{i,1}\\ \alpha_{i,2}\\ \vdots\\ \alpha_{i,N_{\mathsf{0}}}\end{pmatrix}\in\mathbb{R}^{N_{\mathsf{0}}},$ (B.39)
of exponents in (B.20) must satisfy the linear system
$\textbf{S}\,\vec{\alpha}=\vec{s},\quad\textbf{S}\in\mathbb{R}^{N_{\mathsf{r}}\times N_{\mathsf{c}}},\quad\vec{s}\in\mathbb{R}^{N_{\mathsf{r}}},\quad N_{\mathsf{c}}\coloneqq N_{\mathsf{l}}\,N_{\mathsf{0}},\quad N_{\mathsf{r}}\coloneqq N_{\mathsf{l}}\,N_{\mathsf{p}}-\mathrm{rank}(\textbf{Y}),$ (B.40)
where the entries of S and $\vec{s}$ are explicitly known real numbers identified in (B.45)-(B.46) and the components of the matrix $\textbf{Y}\in\mathbb{R}^{(N_{\mathsf{l}}\,N_{\mathsf{p}})\times(N_{\mathsf{p}}\,N_{\mathsf{x}})}$ are explicitly known, as specified by (B.38).

### v2

@@ Appendix D Proof of Theorem 4
We consider the vectors $\vec{l}_{1},\dots,\vec{l}_{N_{\mathsf{l}}}\in\mathbb{R}^{N_{\mathsf{p}}}$ of exponents in (C.1). As explained in the proof of Theorem 3 (Appendix C), we have that Theorems 1 and 2 combined with Assumptions˜4, 5 and 6 ensure (C.2) for each vector $\vec{l}_{1},\dots,\vec{l}_{N_{\mathsf{l}}}$. This can be understood as follows. Let us denote B as the matrix whose columns are the vectors $\vec{b}_{1},\dots,\vec{b}_{N_{\mathsf{0}}}$ (21), i.e.,
@@
$\textbf{B}\coloneqq\begin{pmatrix}\vec{b}_{1}&\vec{b}_{2}&\cdots&\vec{b}_{N_{\mathsf{0}}}\end{pmatrix}\in\mathbb{R}^{N_{\mathsf{p}}\times N_{\mathsf{0}}},\quad\mbox{with}\quad N_{\mathsf{0}}<N_{\mathsf{p}}.$ (D.1)
@@
Each of the vectors $\vec{l}_{1},\dots,\vec{l}_{N_{\mathsf{l}}}$ in (C.2) belongs to the column space of such a matrix B (D.1), which is equivalent to the orthogonal complement of the null space (kernel) of $\textbf{B}^{T}$, i.e.,
@@
$\vec{l}_{i}\in\mathrm{colsp}(\textbf{B})\equiv\left(\mathrm{ker}(\textbf{B}^{T})\right)^{\bot},\quad\forall i=1,\dots,N_{\mathsf{l}}.$ (D.2)
@@
One has $N_{\mathsf{0}}=N_{\mathsf{p}}-\mathrm{rank}(\textbf{M})<N_{\mathsf{p}}$ in (D.1) owing to $\mathrm{rank}(\textbf{M})\geq 1$. The latter inequality follows from Assumption˜6 guaranteeing that M (20) cannot be a matrix whose entries are all equal to zero. Then, since $\mathrm{rank}(\textbf{B}^{T})=\mathrm{rank}(\textbf{B})=N_{\mathsf{0}}$ and $\mathrm{dim}\left(\mathrm{ker}(\textbf{B}^{T})\right)=N_{\mathsf{p}}-N_{\mathsf{0}}>0$, we can consider $N_{\mathsf{p}}-N_{\mathsf{0}}$ vectors $\vec{v}_{1},\dots,\vec{v}_{N_{\mathsf{p}}-N_{\mathsf{0}}}\in\mathbb{R}^{N_{\mathsf{p}}}$ to form a basis of $\mathrm{ker}(\textbf{B}^{T})$. The assertion (D.2) is equivalently stated as each of $\vec{l}_{1},\dots,\vec{l}_{N_{\mathsf{l}}}$ being perpendicular to each of the basis’ vectors $\vec{v}_{1},\dots,\vec{v}_{N_{\mathsf{p}}-N_{\mathsf{0}}}$, i.e.,
@@
$\textbf{V}^{T}\,\vec{l}_{i}=\vec{0},\quad\textbf{V}\coloneqq\begin{pmatrix}\vec{v}_{1}&\vec{v}_{2}&\cdots&\vec{v}_{N_{\mathsf{p}}-N_{\mathsf{0}}}\end{pmatrix}\in\mathbb{R}^{N_{\mathsf{p}}\times(N_{\mathsf{p}}-N_{\mathsf{0}})},\quad\forall i=1,\dots,N_{\mathsf{l}}.$ (D.3)
@@
Assumption˜5 guarantees that the numerical values $\kappa_{i}$ and $\theta_{j}$ of all $\tilde{\kappa}_{i}$ in (15) and $\tilde{\theta}_{j}$ in (16) are strictly positive. Then, one can plug (15) and (16) into (14) to get
$\lambda_{i}=\tilde{p}_{1}^{\sum_{j=1}^{N_{\mathsf{x}}}c_{i,j}\,t_{1,j}\,+\,d_{i,1}}\,\tilde{p}_{2}^{\sum_{j=1}^{N_{\mathsf{x}}}c_{i,j}\,t_{2,j}\,+\,d_{i,2}}\,\cdots\,\tilde{p}_{N_{\mathsf{p}}}^{\sum_{j=1}^{N_{\mathsf{x}}}c_{i,j}\,t_{N_{\mathsf{p}},j}\,+\,d_{i,N_{\mathsf{p}}}},\quad\forall i=1,\dots,N_{\mathsf{l}}.$ (D.4)
@@
By using (D.4) and the independence of physical parameters $\tilde{p}_{1},\dots,\tilde{p}_{N_{\mathsf{p}}}$ (Assumption˜4), the vectors $\vec{l}_{1},\dots,\vec{l}_{N_{\mathsf{l}}}$ of exponents in (C.1) can be computed in terms of $\vec{t}_{1},\dots,\vec{t}_{N_{\mathsf{x}}}$ (23):
$\vec{l}_{i}=c_{i,1}\,\vec{t}_{1}\,+\,c_{i,2}\,\vec{t}_{2}\,+\,\cdots\,+\,c_{i,N_{\mathsf{x}}}\,\vec{t}_{N_{\mathsf{x}}}\,+\,\vec{d}_{i},\quad\vec{d}_{i}\coloneqq(d_{i,1},d_{i,2},\dots,d_{i,{N_{\mathsf{p}}}})^{T},\quad\forall i=1,\dots,N_{\mathsf{l}}.$ (D.5)
Inserting $\vec{l}_{i}$ (D.5) into (D.3) gives
$c_{i,1}\,\textbf{V}^{T}\,\vec{t}_{1}\,+\,c_{i,2}\,\textbf{V}^{T}\,\vec{t}_{2}\,+\,\cdots\,+\,c_{i,N_{\mathsf{x}}}\,\textbf{V}^{T}\,\vec{t}_{N_{\mathsf{x}}}=-\textbf{V}^{T}\,\vec{d}_{i},\quad\forall i=1,\dots,N_{\mathsf{l}},$ (D.6)
that reads in terms of the vector $\vec{t}$ (23) of unknowns as
$\textbf{T}\,\vec{t}=\vec{\tau},\quad\textbf{T}\in\mathbb{R}^{N_{\mathsf{t}}\times(N_{\mathsf{p}}\,N_{\mathsf{x}})},\quad\vec{\tau}\in\mathbb{R}^{N_{\mathsf{t}}},\quad N_{\mathsf{t}}\coloneqq N_{\mathsf{l}}\,(N_{\mathsf{p}}-N_{\mathsf{0}}),$ (D.7)
@@
$\textbf{T}=\begin{pmatrix}c_{1,1}\,\textbf{V}^{T}&c_{1,2}\,\textbf{V}^{T}&\cdots&c_{1,N_{\mathsf{x}}}\,\textbf{V}^{T}\\ c_{2,1}\,\textbf{V}^{T}&c_{2,2}\,\textbf{V}^{T}&\cdots&c_{2,N_{\mathsf{x}}}\,\textbf{V}^{T}\\ \vdots&\vdots&\ddots&\vdots\\ c_{N_{\mathsf{l}},1}\,\textbf{V}^{T}&c_{N_{\mathsf{l}},2}\,\textbf{V}^{T}&\cdots&c_{N_{\mathsf{l}},N_{\mathsf{x}}}\,\textbf{V}^{T}\end{pmatrix},\quad\vec{\tau}=-\begin{pmatrix}\textbf{V}^{T}\,\vec{d}_{1}\\ \textbf{V}^{T}\,\vec{d}_{2}\\ \vdots\\ \textbf{V}^{T}\,\vec{d}_{N_{\mathsf{l}}}\end{pmatrix}$, (D.8)
@@
with V and $\vec{d}_{1},\dots,\vec{d}_{N_{\mathsf{l}}}$ defined as in (D.3) and (D.5), respectively. By using any solution $\vec{t}\coloneqq(\vec{t}_{1},\vec{t}_{2},\dots,\vec{t}_{N_{\mathsf{x}}})^{T}$ to (D.7)-(D.8) in (D.5), the resulting vectors $\vec{l}_{1},\dots,\vec{l}_{N_{\mathsf{l}}}$ (D.5) fulfill (D.3) that implies (D.2). In other words, such $\vec{l}_{1},\dots,\vec{l}_{N_{\mathsf{l}}}$ belong to the column space of the matrix B (D.1) generated by the basis’ vectors $\vec{b}_{1},\dots,\vec{b}_{N_{\mathsf{0}}}$ (21). Thus, each of the computed $\vec{l}_{1},\dots,\vec{l}_{N_{\mathsf{l}}}$ admits the unique vector $\vec{\alpha}_{i}$ (C.2) of coordinates with respect to the basis $\vec{b}_{1},\dots,\vec{b}_{N_{\mathsf{0}}}$ (21). That is to say, each of the $N_{\mathsf{l}}$ linear systems written in (C.2), i.e., $\textbf{B}\,\vec{\alpha}_{i}=\vec{l}_{i}$, with $i=1,\dots,N_{\mathsf{l}}$, can be solved providing the unique solution $\vec{\alpha}_{i}\in\mathbb{R}^{N_{\mathsf{0}}}$:
@@
$\vec{\alpha}_{i}=\textbf{B}^{+}\,\vec{l}_{i},\quad\textbf{B}^{+}\coloneqq\left(\textbf{B}^{T}\,\textbf{B}\,\right)^{-1}\textbf{B}^{T}\in\mathbb{R}^{N_{\mathsf{0}}\times N_{\mathsf{p}}},\quad\forall i=1,\dots,N_{\mathsf{l}},$ (D.9)
@@
where $\textbf{B}^{+}$ is the pseudo-inverse of the rectangular matrix B (D.1). The inverse matrix $\left(\textbf{B}^{T}\,\textbf{B}\,\right)^{-1}\in\mathbb{R}^{N_{\mathsf{0}}\times N_{\mathsf{0}}}$ in (D.9) is well defined because $\mathrm{rank}(\textbf{B})=N_{\mathsf{0}}<N_{\mathsf{p}}$ implies that $\mathrm{rank}(\textbf{B}^{T}\,\textbf{B})=N_{\mathsf{0}}$. By plugging $\vec{l}_{i}$ (D.5) into (D.9), it follows
$c_{i,1}\,\textbf{B}^{+}\,\vec{t}_{1}\,+\,c_{i,2}\,\textbf{B}^{+}\,\vec{t}_{2}\,+\,\cdots\,+\,c_{i,N_{\mathsf{x}}}\,\textbf{B}^{+}\,\vec{t}_{N_{\mathsf{x}}}=\vec{\alpha}_{i}\,-\,\textbf{B}^{+}\,\vec{d}_{i},\quad\forall i=1,\dots,N_{\mathsf{l}},$ (D.10)
that reads in terms of the vector $\vec{t}$ (23) of unknowns as
$\textbf{A}\,\vec{t}=\vec{\alpha}\,-\,\vec{\omega},\quad\textbf{A}\in\mathbb{R}^{(N_{\mathsf{l}}\,N_{\mathsf{0}})\times(N_{\mathsf{p}}\,N_{\mathsf{x}})},\quad\vec{\alpha},\vec{\omega}\in\mathbb{R}^{N_{\mathsf{l}}\,N_{\mathsf{0}}},$ (D.11)
@@
$\textbf{A}=\begin{pmatrix}c_{1,1}\,\textbf{B}^{+}&c_{1,2}\,\textbf{B}^{+}&\cdots&c_{1,N_{\mathsf{x}}}\,\textbf{B}^{+}\\ c_{2,1}\,\textbf{B}^{+}&c_{2,2}\,\textbf{B}^{+}&\cdots&c_{2,N_{\mathsf{x}}}\,\textbf{B}^{+}\\ \vdots&\vdots&\ddots&\vdots\\ c_{N_{\mathsf{l}},1}\,\textbf{B}^{+}&c_{N_{\mathsf{l}},2}\,\textbf{B}^{+}&\cdots&c_{N_{\mathsf{l}},N_{\mathsf{x}}}\,\textbf{B}^{+}\end{pmatrix},\quad\vec{\omega}=\begin{pmatrix}\textbf{B}^{+}\,\vec{d}_{1}\\ \textbf{B}^{+}\,\vec{d}_{2}\\ \vdots\\ \textbf{B}^{+}\,\vec{d}_{N_{\mathsf{l}}}\end{pmatrix}$, (D.12)
@@
where $\vec{\alpha}$ is defined as in (25), being the vector $\in\mathbb{R}^{N_{\mathsf{l}}\,N_{\mathsf{0}}}$ whose entries are the exponents $\alpha_{i,k}$ in (22). The matrix $\textbf{B}^{+}$ and the vectors $\vec{d}_{1},\dots,\vec{d}_{N_{\mathsf{l}}}$ are defined as in (D.9) and (D.5), respectively. Both linear systems (D.7)-(D.8) and (D.11)-(D.12) must be satisfied by the vector $\vec{t}$ (23) of unknowns, yielding to the matrix equation $\textbf{Y}\,\vec{t}=\vec{\gamma}$ (24) for $\vec{t}$, with Y and $\vec{\gamma}$ given by
@@
$\textbf{Y}=\begin{pmatrix}\textbf{A}\\ \textbf{T}\end{pmatrix}\in\mathbb{R}^{(N_{\mathsf{l}}\,N_{\mathsf{p}})\times(N_{\mathsf{p}}\,N_{\mathsf{x}})},\quad\vec{\gamma}=\begin{pmatrix}\vec{\alpha}-\vec{\omega}\\ \vec{\tau}\end{pmatrix}\in\mathbb{R}^{N_{\mathsf{l}}\,N_{\mathsf{p}}}.$ (D.13)
@@
The entries of Y (D.13) are explicitly known real numbers, as given by those of A (D.12) and T (D.8). Since the components of $\vec{\omega}$ (D.12) and $\vec{\tau}$ (D.8) are explicitly known, the entries of $\vec{\gamma}$ (D.13) are explicit affine functions of the exponents $\alpha_{i,k}$ in (22), collected by the vector $\vec{\alpha}$ (25).
Appendix E Proof of Theorem 5
See 5